Describe a pattern in the sequence of numbers. Predict the next number.
1) 1, 2, 6, 24, ___
2) 0, 3, 8, 15, 24, ___

Found 2 solutions by sudhanshu_kmr, Edwin McCravy:
Answer by sudhanshu_kmr(1152)

1) 1*2= 2
2*3= 6
6*4= 24
24*5= 120
next number is 120.

2) 0+3 = 3
3+5= 8
8+7=15
15+9=24
24+11=35
next number is 35.

Answer by Edwin McCravy(20059)

1) 1, 2, 6, 24, ___
2) 0, 3, 8, 15, 24, ___
List the numbers this way

n  term
1    1
2    2
3    6
4   24 
5

Now I'll do some coloring:

n term
1   1 
2   2
3   6
4  24 
5

We notice above that if we multiply the red 1 by the blue 2 we get the green 2.

n term
1   1
2   2
3   6
4  24 
5

We notice above that if we multiply the red 2 by the blue 3 we get the green 6.

 
n term
1   1
2   2
3   6
4  24
5

We notice above that if we multiply the red 6 by the blue 4 we get the green 24.


n term
1   1
2   2
3   6
4  24 
5 120

We REASON that in the above if we multiply the red 24 by the blue 5 we must get a green 120. 

[This sequence is the sequence of "factorials", indicated by n!, but you don't need
 to know that yet].

-----------------

2) 0, 3, 8, 15, 24, ___

n term
1   0
2   3
3   8
4  15 
5  24
6

Think of another sequence that this one is near.  How about this sequence:

n term
1   1
2   4
3   9
4  16 
5  25
6  36

Each term of our sequence is 1 less than this sequence and this sequence is easily
recognized as the sequence of the squares of n or nē.

So every term of your sequence is 1 less than the corresponding sequence
of squares,

so the next term has to be 36-1 or 35.

Edwin